A violin string vibrating in the 1st harmonic produces a note of frequency 316 Hz. If the speed of the wave energy along the violin string is 280 m/s, calculate the length of the string.


1
Expert's answer
2021-01-30T06:25:41-0500

For the 1st harmonic


"\nu=\frac{v}{2L}\to L=\frac{v}{2\nu}=\frac{280}{2\cdot316}=0.443(m)" . Answer